# Analytical model of precessing binaries using post-Newtonian theory in the extreme mass-ratio limit I: General Formalism

### 14 Feb 2024 v1

@@ I Introduction
Projections based on future LISA detections show that tests of strong gravity with EMRIs will improve current constraints by several orders of magnitude Babak et al. (2017); Barausse et al. (2020); Arun et al. (2022); Baibhav et al. (2021), including searches for extra fundamental charges and fields Cardoso et al. (2011); Yunes et al. (2012); Pani et al. (2011); Maselli et al. (2020, 2022); Barsanti et al. (2022b, 2023); Liang et al. (2023); Zhang et al. (2023); Zi et al. (2023); Lestingi et al. (2023), anomalous multipole moments Barack and Cutler (2007); Babak et al. (2017); Fransen and Mayerson (2022); Raposo et al. (2019); Bena and Mayerson (2020); Bianchi et al. (2020); Loutrel et al. (2022), tests of the Kerr bound on the spin of the secondary Piovano et al. (2020a), nonvanishing tidal Love numbers of the primary Pani and Maselli (2019); Piovano et al. (2023), horizon-scale structure Datta et al. (2020); Datta and Bose (2019); Maggio et al. (2021), tests of exotic compact objects Pani et al. (2010); Macedo et al. (2013); Destounis et al. (2023) and of ultralight-boson condensates around BHs Hannuksela et al. (2019, 2020); Brito and Shah (2023); Duque et al. (2023).
@@ IV.5.2 $\delta\Phi$ at ${\cal{O}}(q)$
where $[{\cal{B}}_{1}^{C},{\cal{B}}_{1}^{S}]$ are functions of $\bar{\alpha}$ given explicitly in Appendix C. We stop our computation here, since it suffices for capturing the leading order behavior of the secondary spin. The methodology can be extended to higher order is one desires more phase accuracy.
@@ V Waveform Phasing Due To Spin Precession And Secondary Spin
where $f_{\rm ISCO}$ is the Fourier frequency associated with the Kerr ISCO, and $f_{\rm 1yr}$ is the Fourier frequency one year before the EMRI reaches the ISCO. Then, the de-phasing between an EMRI with a spinning secondary and one without is
@@
where $\theta^{a}$ is held fixed between the two inspirals. The above quantity is useful to study, since, assuming the absence of parameter degeneracies, a de-phasing greater than one radian would substantially impact a matched-filter search, leading to a significant loss of detected events, and being potentially detectable Lindblom et al. (2008).
@@
For this analysis, we use the TaylorF2 approximation for $\Psi(f)$ given in Eq. (IV.3), while for $\delta\Phi(f)$ we use the analytic approximation given in Sec. IV.5 with $\alpha$ and $\gamma_{P}$ given by Eqs. (131) & (132), respectively. The quantity $\delta\Psi_{\chi_{2}}$ is dependent on the parameters $\theta^{a}$, but many of these have negligible impact on the de-phasing. The three most important parameters we have identified are $q$, $L_{z,0}$, and $\chi_{1}$. The first should be obvious since the secondary’s spin is coupled to the mass ratio. The z-component of the orbital angular momentum’s initial orientation $L_{z,0}$ controls the amplitude of precession effects, specifically, more misalignment (smaller $L_{z,0}$) leads to stronger precession effects. Lastly, the primary’s spin controls the location of the ISCO, and the total angular momentum through Eq. (110), modulating how much the primary influences the precession of the secondary. For the analysis carried out here, we fix the orientation of the line of sight vector $\vec{N}$, which enters the precession phase $\delta\Phi$, to $\theta_{N}=\pi/6$ and $\phi_{N}=\pi/4$.
@@
Fig. 5 shows the de-phasing $\delta\Psi_{\chi_{2}}$ as a function of frequency (left) and time (right) for various values of the primary’s spin, and with $\bar{\chi}_{Q,0}=1$ and $\bar{\chi}_{P,0}=0=\dot{\bar{\chi}}_{P,0}$. In this case, the secondary’s spin is a constant in the co-precessing reference frame of Sec. III.2, and thus only undergoes simple precession (no nutation). The total de-phasing over one year of inspiral is typically $0.7-1.0$ radian, depending on the value of the primary’s spin $\chi_{1}$. Larger values of of $\chi_{1}$ produce slightly more de-phasing than lower values, owing to the ISCO being smaller for prograde orbits and allowing for a larger number of total GW cycles. Fig. 6 shows the same comparison, but for the secondary spin with orientation $[\bar{\chi}_{Q,0},\bar{\chi}_{P,0},\dot{\bar{\chi}}_{P,0}]=[0.687,-0.259,-0.635]$. The de-phasings in this case are slightly smaller than those in Fig. 5, due to the fact that the components $[\bar{\chi}_{P,0},\dot{\bar{\chi}}_{P,0}]$ mainly contribute to the phase through oscillatory corrections, which are suppressed by $q^{2}$. Thus, the component $\bar{\chi}_{Q,0}$ produces the largest contribution to the phase, and will likely be the best recovered component of the secondary’s spin when performing parameter estimation.
@@
In Fig. 7, we investigate how the initial misalignment of the orbital angular momentum, encoded in $L_{z,0}$, impacts the phase accumulated due to the secondary’s spin. The limit $L_{z,0}\rightarrow 1$ corresponds to alignment between the orbital angular momentum and primary spin, and decreasing values provide greater misalignment. Generally, more misalignment produces more precession cycles, thus increasing the impact of the secondary spin on the waveform’s phase. For the largest misalignment studied here, specifically $L_{z,0}=0.1$, $\delta\Psi_{\chi_{2}}\sim O(10)$ radians after one year of inspiral. The case $L_{z,0}=0.87$ provides the smallest de-phasing, which is due to the fact that $\hat{L}$ is nearly aligned with the line of sight vector $\vec{N}$. Lastly, in Fig. 8, we study the impact of the mass ratio on the dephasing. Due to the fact that spin effects are suppressed by the mass ratio, more comparable mass systems (larger $q$) produce more dephasing over one year of inspiral than more disparate mass systems.
@@
Figure 7: Same as Fig. (5) but for $\chi_{1}=0.90$ and varying the value of $L_{z,0}$. Smaller values of $L_{z,0}$ correspond to systems with greater intial misalignment between the orbital angular momentum and primary spin, leading to more precession cycles in the waveform, and thus, larger $\delta\Psi_{\chi_{2}}$.
@@ Acknowledgements.
This work is partially supported by the MIUR PRIN Grant 2020KR4KN2 “String Theory as a bridge between Gauge Theories and Quantum Gravity”, by the FARE programme (GW-NEXT, CUP: B84I20000100001), and by the EU Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie Grant Agreement No. 101007855. N.L. is supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, and the ICSC National Research Centre funded by NextGenerationEU. S.M. is thankful to the GravNet grant for providing financial support to carry out the present research, and Sapienza University of Rome for a warm hospitality where a part of this work was done. S.M. is also thankful to the Inspire Faculty Grant DST/INSPIRE/04/2022/001332, DST, Government of India, and Lumina Quaeruntur No. LQ100032102 of the Czech Academy of Sciences, for support.

### 28 Feb 2024 v2

@@ I Introduction
Projections based on future LISA detections show that tests of strong gravity with EMRIs will improve current constraints by several orders of magnitude Babak et al. (2017); Barausse et al. (2020); Arun et al. (2022); Baibhav et al. (2021), including searches for extra fundamental charges and fields Cardoso et al. (2011); Yunes et al. (2012); Pani et al. (2011); Maselli et al. (2020, 2022); Barsanti et al. (2022b, 2023); Liang et al. (2023); Zhang et al. (2023); Zi et al. (2023); Lestingi et al. (2023), anomalous multipole moments Barack and Cutler (2007); Babak et al. (2017); Fransen and Mayerson (2022); Raposo et al. (2019); Bena and Mayerson (2020); Bianchi et al. (2020); Loutrel et al. (2022); Kumar et al. (2023), tests of the Kerr bound on the spin of the secondary Piovano et al. (2020a), nonvanishing tidal Love numbers of the primary Pani and Maselli (2019); Piovano et al. (2023), horizon-scale structure Datta et al. (2020); Datta and Bose (2019); Maggio et al. (2021), tests of exotic compact objects Pani et al. (2010); Macedo et al. (2013); Destounis et al. (2023) and of ultralight-boson condensates around BHs Hannuksela et al. (2019, 2020); Brito and Shah (2023); Duque et al. (2023).
@@ IV.5.2 $\delta\Phi$ at ${\cal{O}}(q)$
where $[{\cal{B}}_{1,2}^{C},{\cal{B}}_{1,2}^{S}]$ are functions of $\bar{\alpha}$ given explicitly in Appendix C. We stop our computation here, since it suffices for capturing the leading order behavior of the secondary spin. The methodology can be extended to higher order is one desires more phase accuracy.
@@ V Waveform Phasing Due To Spin Precession And Secondary Spin
where $f_{\rm ISCO}$ is the Fourier frequency associated with the Kerr ISCO, and $f_{\rm 1yr}$ is the Fourier frequency one year before the EMRI reaches the ISCO. From this, we define two quantities. The first is the de-phasing between an EMRI with a spinning secondary and one without is
@@
where $\theta^{a}$ is held fixed between the two inspirals. The above quantity is useful to study, since, assuming the absence of parameter degeneracies, a de-phasing greater than one radian would substantially impact a matched-filter search, leading to a significant loss of detected events, and being potentially detectable Lindblom et al. (2008). The second quantity is the de-phasing between an aligned (non-precessing) EMRI and an EMRI undergoing spin-orbit precession (without a spinning secondary), specifically
$\delta\Psi_{L}(\theta^{a})=\Delta\Psi_{T}(\theta^{a},\chi_{2}^{a}=0)-\Delta\Psi_{T}(\theta^{a}_{\rm align},\chi_{2}^{a}=0)\,,$ (161)
where $\theta^{a}_{\rm align}$ are the parameters in the aligned limit.
To begin with, we consider how many radians of phase are introduced by the effect of spin-orbit precession, specifically $\delta\Psi_{L}$ in Eq. (161). For this comparison, the spin of the secondary is neglected, while we take $\chi_{\rm eff}=\chi_{1}L_{z,0}$, which follows from the expansion of Eq. (11) about $q\ll 1$. The aligned limit, specified by $\theta^{a}_{\rm align}$, is given by $L_{z,0}=1$. We use the TaylorF2 approximation for $\Psi(f)$ given in Eq. (IV.3), while for $\delta\Phi(f)$ we use the analytic approximation given in Sec. IV.5 with $\alpha$ and $\gamma_{P}$ given by Eqs. (131) & (132), respectively. For the analysis carried out here, we fix the orientation of the line of sight vector $\vec{N}$, which enters the precession phase $\delta\Phi$, to $\theta_{N}=\pi/6$ and $\phi_{N}=\pi/4$. We fix $\chi_{1}=0.9$ and $q=10^{-5}$, and study how the de-phasing $\delta\Psi_{L}$ varies with increasing misalignment, i.e. with decreasing $L_{z,0}$. Fig. 5 shows the results of this comparison, revealing that increasing misalignment produces greater de-phasing compared to the aligned limit, owing to the increasing precession effects on the binary. The total dephasing is generally large, typically $\gtrsim 10^{4}$ radians or larger, even for EMRIs with small misalignment. The reason for this is that the misalignment enters the GW phase at leading order in the mass ratio ${\cal{O}}(q^{-1})$ and at 1.5PN order, ${\cal{O}}(v^{3})$, through $\chi_{\rm eff}$. Thus, the contribution for generic spin-orbit precession is large.
@@
We now consider the contribution of a spinning secondary to the GW phase. The quantity $\delta\Psi_{\chi_{2}}$ is dependent on the parameters $\theta^{a}$, but many of these have negligible impact on the de-phasing. The three most important parameters we have identified are $q$, $L_{z,0}$, and $\chi_{1}$. The first should be obvious since the secondary’s spin is coupled to the mass ratio. The z-component of the orbital angular momentum’s initial orientation $L_{z,0}$ controls the amplitude of precession effects, specifically, more misalignment (smaller $L_{z,0}$) leads to stronger precession effects. Lastly, the primary’s spin controls the location of the ISCO, and the total angular momentum through Eq. (110), modulating how much the primary influences the precession of the secondary.
@@
Fig. 6 shows the de-phasing $\delta\Psi_{\chi_{2}}$ as a function of frequency (left) and time (right) for various values of the primary’s spin, and with $\bar{\chi}_{Q,0}=1$ and $\bar{\chi}_{P,0}=0=\dot{\bar{\chi}}_{P,0}$. In this case, the secondary’s spin is a constant in the co-precessing reference frame of Sec. III.2, and thus only undergoes simple precession (no nutation). The total de-phasing over one year of inspiral is typically $0.7-1.0$ radian, depending on the value of the primary’s spin $\chi_{1}$. Larger values of of $\chi_{1}$ produce slightly more de-phasing than lower values, owing to the ISCO being smaller for prograde orbits and allowing for a larger number of total GW cycles. Fig. 7 shows the same comparison, but for the secondary spin with orientation $[\bar{\chi}_{Q,0},\bar{\chi}_{P,0},\dot{\bar{\chi}}_{P,0}]=[0.687,-0.259,-0.635]$. The de-phasings in this case are slightly smaller than those in Fig. 6, due to the fact that the components $[\bar{\chi}_{P,0},\dot{\bar{\chi}}_{P,0}]$ mainly contribute to the phase through oscillatory corrections, which are suppressed by $q^{2}$. Thus, the component $\bar{\chi}_{Q,0}$ produces the largest contribution to the phase, and will likely be the best recovered component of the secondary’s spin when performing parameter estimation. Note that this is not unexpected, since it is known that the component of the secondary spin parallel to the total angular momentum, which is related to $\bar{\chi}_{Q,0}$, produces the dominate contribution to the GW phase Witzany (2019); Witzany and Piovano (2023); Skoupy et al. (2023); Burke et al. (2023).
@@
In Fig. 8, we investigate how the initial misalignment of the orbital angular momentum, encoded in $L_{z,0}$, impacts the phase accumulated due to the secondary’s spin. The limit $L_{z,0}\rightarrow 1$ corresponds to alignment between the orbital angular momentum and primary spin, and decreasing values provide greater misalignment. Generally, more misalignment produces more precession cycles, thus increasing the impact of the secondary spin on the waveform’s phase. For the largest misalignment studied here, specifically $L_{z,0}=0.1$, $\delta\Psi_{\chi_{2}}\sim{\cal{O}}(10)$ radians after one year of inspiral. The case $L_{z,0}=0.87$ provides the smallest de-phasing, which is due to the fact that $\hat{L}$ is nearly aligned with the line of sight vector $\vec{N}$. Lastly, in Fig. 9, we study the impact of the mass ratio on the dephasing. Due to the fact that spin effects are suppressed by the mass ratio, more comparable mass systems (larger $q$) produce more dephasing over one year of inspiral than more disparate mass systems.
@@
Figure 5: GW de-phasing (161) between an aligned EMRI $(L_{z,0}=1)$ and a precessing configuration with varying misalignment (colored curves, smaller values of $L_{z,0}$ correspond to large misalignment). The EMRIs in this figure has a non-spinning secondary, and the precession is induced by the spin-orbit coupling between the primary spin and orbital angular momentum. Further, we fix $\chi_{1}=0.90$ and $q=10^{-5}$.
@@
Figure 8: Same as Fig. (6) but for $\chi_{1}=0.90$ and varying the value of $L_{z,0}$. Smaller values of $L_{z,0}$ correspond to systems with greater initial misalignment between the orbital angular momentum and primary spin, leading to more precession cycles in the waveform, and thus, larger $\delta\Psi_{\chi_{2}}$.
@@ Acknowledgements.
We thank Gabriel Piovano for interesting comments on the draft. This work is partially supported by the MIUR PRIN Grant 2020KR4KN2 “String Theory as a bridge between Gauge Theories and Quantum Gravity”, by the FARE programme (GW-NEXT, CUP: B84I20000100001), and by the EU Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie Grant Agreement No. 101007855. N.L. is supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, and the ICSC National Research Centre funded by NextGenerationEU. S.M. is thankful to the GravNet grant for providing financial support to carry out the present research, and Sapienza University of Rome for a warm hospitality where a part of this work was done. S.M. is also thankful to the Inspire Faculty Grant DST/INSPIRE/04/2022/001332, DST, Government of India, and Lumina Quaeruntur No. LQ100032102 of the Czech Academy of Sciences, for support.
@@ Appendix C Coefficients of the Waveform Precession Phase
For the oscillatory correction in Eq. (IV.5.2), the functions ${\cal{B}}_{1,2}^{C,S}(\bar{\alpha})$ can be decomposed as
${\cal{B}}_{1}^{C,S}$ $=\nu_{SL}\sum_{k=0}^{2}\left[{\cal{H}}^{C,S}_{(k)}\cos(k\bar{\alpha})+{\cal{K}}^{C,S}_{(k)}\sin(k\bar{\alpha})\right]\,,$ (242)
${\cal{B}}_{2}^{C,S}$ $=\frac{v\omega_{SL}\omega_{L}}{\omega_{P}}{\cal{I}}(\bar{\alpha}){\cal{J}}^{C,S}(\alpha)\,,$ (243)
with
${\cal{H}}_{(0)}^{C}$ $=\frac{1}{2}\hat{L}_{z,0}^{2}\nu_{QP}\bar{\chi}_{P,0}\left(\sin^{2}\theta_{N}-2\cos^{2}\theta_{N}\right)\,,$ (244)
${\cal{H}}_{(1)}^{C}$ $=\frac{\hat{L}_{z,0}(2\hat{L}_{z,0}^{2}-1)}{\sqrt{1-\hat{L}_{z,0}^{2}}}\nu_{QP}\bar{\chi}_{P,0}\cos\theta_{N}\sin\theta_{N}\,,$ (245)
${\cal{H}}_{(2)}^{C}$ $=\frac{1}{2}\hat{L}_{z,0}^{2}\nu_{QP}\bar{\chi}_{P,0}\sin^{2}\theta_{N}\,,$ (246)
${\cal{K}}_{(1)}^{C}$ $=\frac{\hat{L}_{z,0}}{\sqrt{1-\hat{L}_{z,0}^{2}}}\dot{\bar{\chi}}_{P,0}\cos\theta_{N}\sin\theta_{N}\,,$ (247)
${\cal{K}}_{(2)}^{C}$ $=\frac{1}{2}\dot{\bar{\chi}}_{P,0}\sin^{2}\theta_{N}\,,$ (248)
${\cal{H}}_{(0)}^{S}$ $=\frac{1}{2}\hat{L}_{z,0}^{2}\nu_{QP}\dot{\bar{\chi}}_{P,0}\left(\sin^{2}\theta_{N}-2\cos^{2}\theta_{N}\right)$ (249)
${\cal{H}}_{(1)}^{S}$ $=\frac{\hat{L}_{z,0}(2\hat{L}_{z,0}^{2}-1)}{\sqrt{1-\hat{L}_{z,0}^{2}}}\nu_{QP}\dot{\bar{\chi}}_{P,0}\cos\theta_{N}\sin\theta_{N}\,,$ (250)
${\cal{H}}_{(2)}^{S}$ $=\frac{1}{2}\hat{L}_{z,0}^{2}\nu_{QP}\dot{\bar{\chi}}_{P,0}\sin^{2}\theta_{N}\,,$ (251)
${\cal{K}}_{(1)}^{S}$ $=-\frac{\hat{L}_{z,0}}{\sqrt{1-\hat{L}_{z,0}^{2}}}\bar{\chi}_{P,0}\cos\theta_{N}\sin\theta_{N}\,,$ (252)
${\cal{K}}_{(2)}^{S}$ $=-\frac{1}{2}\bar{\chi}_{P,0}\sin^{2}\theta_{N}$ (253)
${\cal{I}}(\bar{\alpha})$ $=\left(1-2\hat{L}_{z,0}^{2}+\hat{L}_{z,0}^{2}\cos^{2}\theta_{N}\right)\cos\theta_{N}$
$-2\hat{L}_{z,0}\sqrt{1-\hat{L}_{z,0}^{2}}\sin^{3}\theta_{N}\cos\bar{\alpha}$
$+\left(1+\hat{L}_{z,0}^{2}\right)\cos\theta_{N}\sin\theta_{N}\cos^{2}\theta_{N}\,,$ (254)
${\cal{J}}^{S}(\alpha)$ $=\dot{\bar{\chi}}_{P,0}\cos\theta_{N}$
$+\left[S_{x}^{(a)}\cos\phi_{N}+S_{y}^{(a)}\sin\phi_{N}\right]\sin\theta_{N}\sin\alpha$
$+\left[C_{x}^{(a)}\cos\phi_{N}+C_{y}^{(a)}\sin\phi_{N}\right]\sin\theta_{N}\cos\alpha\,,$ (255)
${\cal{J}}^{C}(\alpha)$ $=\bar{\chi}_{P,0}\cos\theta_{N}$
$+\left[C_{x}^{(s)}\cos\phi_{N}+C_{y}^{(s)}\sin\phi_{N}\right]\sin\theta_{N}\sin\alpha$
$+\left[S_{x}^{(s)}\cos\phi_{N}+S_{y}^{(s)}\sin\phi_{N}\right]\sin\theta_{N}\cos\alpha\,,$ (256)
$C_{x,y}^{(a)}$ $=C_{x,y}^{(+)}-C_{x,y}^{(-)}\,,\qquad C_{x,y}^{(s)}=C_{x,y}^{(+)}+C_{x,y}^{(-)}\,,$ (257)
$S_{x,y}^{(a)}$ $=S_{x,y}^{(+)}-S_{x,y}^{(-)}\,,\qquad S_{x,y}^{(s)}=S_{x,y}^{(+)}+S_{x,y}^{(-)}\,,$ (258)
where $S_{x,y}^{(\pm)}$ and $C_{x,y}^{(\pm)}$ are given in Eqs. (82)-(83).